OpenStudy (anonymous):

The area of a rectangle is found by multiplying the length by the width: A = lw. A certain rectangle has an area of x2 + 7x + 12. Factor the trinomial to find the length and width of the rectangle. l = x + 6; w = x + 2 l = x + 5; w = x + 7 l = x + 4; w = x + 3

OpenStudy (anonymous):

x^2 + 7x + 12 = x^2 + 3x + 4x + 12 = x(x + 3) + 4(x + 3) = (x + 3)(x + 4)